Factor x^2+3x-4=50
Found 2 solutions by Alan3354, lynnlo:
Answer by Alan3354(69443)

Factor x^2+3x-4=50
---------------
Polynomials are factored.
Equations are solved.
This is an equation.
=======================
Factor x^2+3x-4=50 **** That can't be factored.
----
Subtract 50 from both sides.
x^2 + 3x - 54 = 0
Now the left side can be factored.
(x+9)*(x-6) = 0
--> x = -9, x = 6

Answer by lynnlo(4176)

(x-1)(x+4)=50